How many electrons are present in 1.6g of methane?1.6.023×10^222.6.02×10^253.6.02×10^234.3.01×10^23

Grade:11

1 Answers

Yash dhadbale
6 years ago
No of moles = given wt / molecular wt = 1.6/16 = 0.1 No of atom = mole *atomicity* avagardos no =5*0.1*6.022*10^23 =30.110* 10^22